OpenStudy (anonymous):

You select a card at random from the cards that make up the word “replacement”. Without replacing the card, you choose a second card. Find the probability of choosing a vowel and then not a vowel. There is 1 letter for each card.

OpenStudy (anonymous):

what is the probability you get a vowel on the first try?

OpenStudy (anonymous):

1/4?

OpenStudy (anonymous):

how many vowels?

OpenStudy (anonymous):

it is a matter of counting count the number of vowels, then count the total number of letters it is the ratio of vowels to the total

OpenStudy (anonymous):

so it would be 4/11?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now one letter is gone, it is a vowel how many not vowels are left, and how many letters are left?

OpenStudy (anonymous):

we usually call those "consonants "

OpenStudy (anonymous):

3/10, and 7/11?

OpenStudy (anonymous):

hold on one vowel is gone, because you know you picked a vowel first that leaves 10 cards left, and 7 consonants right?

OpenStudy (anonymous):

so the probability that the second card is a consonant if you know the first card is a vowel is \(\frac{7}{10}\)

OpenStudy (anonymous):

you will find your answer by multiplying \[\frac{4}{11}\times \frac{7}{10}\]

OpenStudy (anonymous):

so it would be 14/55

OpenStudy (anonymous):

yes